# A bound-preserving and conservative enriched Galerkin method for elliptic problems

### v1

@@ 3 Positivity-preserving finite elements
$v_{h}=v^{1}_{h}+v^{0}_{h}\qquad\text{ where }v_{h}\in V_{h},\;v^{1}_{h}\in V^{1}_{h},\;v^{0}_{h}\in V^{0}_{h}\,.$ (7)
Next, we denote by $\phi_{1},\ldots,\phi_{N}$ are the standard basis functions (the ”hat” functions) spanning the space $V_{h}^{1}$. Then, given $w_{h}=w_{h}^{1}+w_{h}^{0}\in V_{h}$ we define the truncation
$P^{w_{h}}_{i}(v_{h})=\max\left[a-\underline{w}_{hi},\;\min\left(v^{1}_{h}(\boldsymbol{x}_{i}),b-\overline{w}_{hi}\right)\right],\qquad\text{ where}$ (8)
@@
Additionally, we introduce the operator $Q^{w_{h}}_{h}(v_{h})=v^{1}_{h}-[P^{w_{h}}_{h}(v_{h})]^{1}$, and emphasize that the superscripts 00 and $1$ always refer to the decomposition according to (7). Finally, we abbreviate
@@
where $v^{1+}_{h}=\sum_{i=1}^{N}P^{v_{h}}_{i}(v_{h})\phi_{i}$.
@@
Lemma 1.
Let $v_{h}\in V_{h}$. If $a-\underline{v}_{hi}\leq b-\overline{v}_{hi}$ for all node indices $i=1,\dots,N$, then $v^{+}_{h}(\boldsymbol{x})\in\mathcal{G}$ for all $\boldsymbol{x}\in\Omega$. .
@@ 3.1 The nonlinear finite element method
$s_{h}(w_{h},v_{h})=\alpha\sum_{i=1}^{N}(\epsilon h_{i}^{d-2}+\mu h_{i}^{d})w^{1}_{h}(\boldsymbol{x}_{i})v^{1}_{h}(\boldsymbol{x}_{i}).$ (10)
@@
The finite element method just presented is nonlinear since $u_{h}^{+}$ is a nonlinear transformation of $u_{h}$. So, it requires an appropriate linearization. In addition, as mentioned earlier, the presence of the term $h_{F}^{-\beta}$ in the stabilization can affect the condition number of the linear subproblems tremendously. To address these two issues, we now present a fixed-point iterative algorithm that enables us to both prove the existence of solutions and circumvent the potential ill-conditioning of the scheme. For this, we define the mapping
$\widetilde{T}\colon V^{0}_{h}\ni w^{0}_{h}\mapsto w^{1}_{h}\mapsto\tilde{w}^{0}_{h}\in V^{0}_{h}\,,$ (11a)
@@
Step 1: Compute $w^{1}_{h}\in V_{h}^{1}$ by solving the nonlinear scheme: $a_{h}([P^{w^{0}_{h}}_{h}w_{h}]^{1},v^{1}_{h})+s_{h}([Q^{w^{0}_{h}}_{h}w_{h}]^{1},v^{1}_{h})=(f,v^{1}_{h})_{\mathcal{T}_{h}}-a_{h}(w^{0}_{h},v^{1}_{h})\qquad\forall v^{1}_{h}\in V^{1}_{h}.$ (11b) Step 2: Compute $\tilde{w}^{0}_{h}\in V^{0}_{h}$ as solution of $a_{h}(\tilde{w}^{0}_{h},v^{0}_{h})=(f,v^{0}_{h})_{\mathcal{T}_{h}}-a_{h}(w^{1+}_{h},v^{0}_{h})-\underbrace{s_{h}(w^{1-}_{h},v^{0}_{h})}_{=0}\qquad\forall v^{0}_{h}\in V^{0}_{h}.$ (11c)
@@
    a fixed-point of $\widetilde{T}$ solves (9), which we can easily see by adding (11b) and (11c). So, in the next section we will show that $\widetilde{T}$ has a fixed point;
@@ 4 Existence of a discrete solution
This section establishes the existence and stability of a solution to (9). To this end, we state our main result and prove it in the remainder of this section.
Theorem 1.
@@
$\|[\![u^{0}_{h}]\!]\|^{2}_{\mathcal{F}_{h}}\leq\,C\,\frac{h^{2\beta-4}}{\gamma^{2}}\left[C_{P}^{2}\mu^{-1}\|f\|^{2}_{0,\Omega}+(C_{P}^{2}\mu+\epsilon)\max(|a|,|b|)^{2}|\Omega|\right]\,,$ (12)
@@ 4.1 Preliminaries and auxiliary results
We exploit our fixed-point iteration (11a) to prove this result. Thus, we need to prove that (11b) and (11c) induce well-posed problems, which we perform in Lemmas 6 and 8, respectively. To this end, we need several auxiliary results.
Lemma 2 (Broken Poincaré inequality, [Bre03]).
@@
Lemma 3 (Relation of bilinear forms, (3.8) in [BGPV24]).
@@
$\left(\mu\|v_{h}^{1}\|_{\mathcal{T}_{h}}^{2}+\epsilon\|\nabla v_{h}^{1}\|_{\mathcal{T}_{h}}^{2}\right)^{1/2}=\|v^{1}_{h}\|_{a_{h}}\leq\frac{C_{\textup{equiv}}}{\alpha}\|v^{1}_{h}\|_{s_{h}}\qquad\text{ for all }v^{1}_{h}\in V^{1}_{h}.$
Lemma 4.
Given $v^{0}_{h}\in V^{0}_{h}$, there are constants $C$ and $\tilde{c}$ (independent of $h$ and $v^{0}_{h}$) such that for any $T\in{\mathcal{T}_{h}}$
$\tilde{c}h_{T}^{2}|v^{1+}_{h}|^{2}_{1,T}\leq\|v^{1+}_{h}\|^{2}_{0,T},\quad\text{ and }\quad\|v^{1+}_{h}\|^{2}_{0,\Omega}\leq C\left[\max(|a|,|b|)^{2}|\Omega|+\|v^{0}_{h}\|^{2}_{0,\Omega}\right]\qquad\text{ for all }v^{1}_{h}\in V^{1}_{h}.$
Proof.
The first inequality follows from a standard local inverse inequality (cf. [EG21a, Lemma 12.1]). To show the second inequality, we start by using the mesh regularity and the result proven in [EG21a, Proposition 12.5], and get to
$\|v^{1+}_{h}\|^{2}_{0,\Omega}\leq\,C\sum_{i=1}^{N}h_{i}^{d}\,[v^{1+}_{h}(\boldsymbol{x}_{i})]^{2}\,.$ (13)
Next, the definition of $v^{1+}_{h}(\boldsymbol{x}_{i})$ implies that $|v^{1+}_{h}(\boldsymbol{x}_{i})|\leq\max(|a|,|b|)+\|v^{0}_{h}\|_{\infty,\omega_{i}}$. Using this fact, the inverse inequality, and the mesh regularity we get to
@@
which finishes the proof. ∎
In addition, we also recall the following trace-inverse inequality (see [EG21a, Lemma 12.10]): There exists $C>0$, depending only on the mesh regularity such that for any $T\in{\mathcal{T}_{h}}$ and a facet $F$ of $T$, and any $v_{h}\in V_{h}$ the following holds
$\|v_{h}\|_{0,F}\leq C\,h_{T}^{-1/2}\|v_{h}\|_{0,T}\,.$ (14)
The next two results are instrumental in the proof that the nonlinear problem from Step 1 (cf. (11b)) is well-posed. The first one is an extension of [BGPV24, Lem. 3.1] to the present (more involved) case, but the proof is identical to that of that result, and thus we omit it.
Lemma 5.
@@
Lemma 6.
@@
$[\tilde{T}_{1}v^{1}_{h}-\tilde{T}_{1}r^{1}_{h},v^{1}_{h}-r^{1}_{h}]\geq C_{M}\|v^{1}_{h}-r^{1}_{h}\|^{2}_{a_{h}}\qquad\text{ for all }v^{1}_{h},r^{1}_{h}\in V^{1}_{h}.$ (16)
Proof.
The proof follows exactly the same lines as that of [BGPV24, Th. 3.2] with [BGPV24, Lem. 3.1: (25)] replaced by Lemma 5. ∎
@@ 4.2 Well-posedness of the iteration defined in (11a)
Lemma 7 (Well-posedness of (11b)).
Under the assumptions of Lemma 6, equation (11b) uniquely defines $w^{1}_{h}\in V^{1}_{h}$ for a given $w^{0}_{h}$. Moreover, if $\hat{w}^{0}_{h},\bar{w}^{0}_{h}\in V^{0}_{h}$, and $\hat{w}^{1}_{h},\bar{w}^{1}_{h}\in V^{1}_{h}$ are their images under $\tilde{T}_{1}$, then the following Lipschitz continuity holds
@@
Proof.
The existence and uniqueness of the solution of (11b) are a direct consequence of the strong monotonicity and continuity of $\tilde{T}_{1}$ (see, e.g., [RR06, Theorem 10.49]). To prove the Lipschitz continuity, let $\hat{w}_{h}^{0},\bar{w}_{h}^{0}\in V_{h}^{0}$ and $\hat{w}_{h}^{1},\bar{w}_{h}^{1}\in V_{h}^{1}$ be their images under $\tilde{T}_{1}$. Then, using (16) and the problem (11b) we get to
@@
and the proof is finished using the continuity of $a_{h}(\cdot,\cdot)$. ∎
@@
Lemma 8 (Well-posedness of (11c)).
@@
$a_{h}(w^{0}_{h},v^{0}_{h})=\mu(\tilde{w}^{0}_{h},v^{0}_{h})_{\mathcal{T}_{h}}+\gamma\left\langle\frac{\epsilon+\mu h_{F}^{2}}{h_{F}^{\beta}}[\![\tilde{w}^{0}_{h}]\!],[\![v^{0}_{h}]\!]\right\rangle_{\mathcal{F}_{h}}\qquad\text{ for all}w^{0}_{h},v^{0}_{h}\in V^{0}_{h}.$ (17)
@@
for every $v_{h}^{0}\in V_{h}^{0}$. Taking $v_{h}^{0}=\hat{w}^{0}_{h}-\bar{w}^{0}_{h}$ and using Cauchy-Schwarz’s inequality on the right-hand side above leads to the result. ∎
4.3 Proof of Theorem 1
@@
Proof.
Let $v^{0}_{h}=\tilde{w}^{0}_{h}$ in (11c) and $\tilde{\gamma}=\tfrac{\gamma}{h^{\beta-1}}$. Then, using (17), the ellipticity of $a_{h}(\cdot,\cdot)$ in $V_{h}^{0}$, (11c), Cauchy-Schwarz’s inequality, the local trace-inverse inequality (14), the broken Poincaré inequality from Lemma 2, and a global inverse inequality, we get to
$\|\sqrt{\mu}\tilde{w}^{0}_{h}\|^{2}_{0,\Omega}+\tilde{\gamma}\|[\![\tilde{w}^{0}_{h}]\!]\|^{2}_{\mathcal{F}_{h}}\leq a_{h}(\tilde{w}^{0}_{h},\tilde{w}^{0}_{h})$
@@
The proof then follows from rearranging terms. ∎
Lemma 10.
Let us assume that $\beta\geq 2$. Then, if $\gamma>0$ is sufficiently large, $T\colon B_{1}(0)\to B_{1}(0)$, where $B_{1}(0)\subset V^{0}_{h}$ is the closed unit ball in $(V_{h}^{0},\|[\![\cdot]\!]\|_{\mathcal{F}_{h}})$.
Proof.
Using Lemma 4, Lemma 9, and the broken Poincaré inequality from Lemma 2 we get to
@@
$\leq\,C\,\frac{h^{2\beta-4}}{\gamma^{2}}\left[C_{P}^{2}\mu^{-1}\|f\|^{2}_{0,\Omega}+(C_{P}^{2}\mu+\epsilon)\max(|a|,|b|)^{2}|\Omega|+C_{P}^{2}\frac{C_{p}^{2}\mu+\epsilon}{\epsilon+\mu h^{2}}\|[\![w^{0}_{h}]\!]\|^{2}_{{\mathcal{F}_{h}}}\right]\,.$ (18)
The result is then proved by noticing that $\gamma$ can always be chosen large enough so that the last term in the right-hand side above can be hidden in the left-hand side, and the resulting right-hand side is smaller than $1$. ∎
The proof of Theorem  1 appears as a corollary of the above results.
Proof of Theorem 1.
Since the EG finite element space is finite-dimensional, the operator $\tilde{T}$ is continuous and maps the closed unit ball (a convex compact set) onto itself. Thus, Brouwer’s fixed-point Theorem (see, e.g. [RR06, Theorem 10.41]) states that $\tilde{T}$ has at least one fixed point in $B_{1}(0)$, which is the desired result. ∎

### v2

@@ 3 Positivity-preserving finite elements
(7) $v_{h}=v^{1}_{h}+v^{0}_{h}\qquad\text{ where }v_{h}\in V_{h},\;v^{1}_{h}\in V^{1}_{h},\;v^{0}_{h}\in V^{0}_{h}\,.$
Next, we denote by $\phi_{1},\ldots,\phi_{N}$ are the standard basis functions (the “hat” functions) spanning the space $V_{h}^{1}$. Then, given $w_{h}=w_{h}^{1}+w_{h}^{0}\in V_{h}$ we define the truncation
(8) $P^{w_{h}}_{i}(v_{h})=\max\left[a-\underline{w}_{hi},\;\min\left(v^{1}_{h}(\boldsymbol{x}_{i}),b-\overline{w}_{hi}\right)\right],\qquad\text{ where}$
@@
Additionally, we introduce the operator $Q^{w_{h}}_{h}(v_{h})=v^{1}_{h}-[P^{w_{h}}_{h}(v_{h})]^{1}$, and emphasize that the superscripts 0 and $1$ always refer to the decomposition according to (7). Finally, we abbreviate
@@
where $v^{1+}_{h}=\sum_{i=1}^{N}P^{v_{h}}_{i}(v_{h})\phi_{i}$. The effect of these operators is illustrated in Figure 1 for the one-dimensional case.
$T_{1}$$T_{2}$$T_{3}$$T_{1}$$T_{2}$$T_{3}$a=0a=0b=1b=1
Figure 1: Illustration of $P_{h}^{v_{h}}$ on three elements $T_{1},T_{2},T_{3}$. Left: $v_{h}$ (cyan) and $v^{1}_{h}$ (orange) and $v_{h}^{0}$ (purple). Right: $v_{h}^{+}$ (cyan) and $v_{h}^{1+}$ (orange) with $v_{h}^{-}$ (purple). In both cases, $a=0$ and $b=1$.
@@
Lemma 3.1.
Let $v_{h}\in V_{h}$. If $a-\underline{v}_{hi}\leq b-\overline{v}_{hi}$ for all node indices $i=1,\dots,N$, then $v^{+}_{h}(\boldsymbol{x})\in\mathcal{G}$ for all $\boldsymbol{x}\in\Omega$.
@@ 3.1 The nonlinear finite element method
(10) $s_{h}(w_{h},v_{h})=\alpha\sum_{i=1}^{N}(\epsilon h_{i}^{d-2}+\mu h_{i}^{d})w^{1}_{h}(\boldsymbol{x}_{i})v^{1}_{h}(\boldsymbol{x}_{i}).$
@@
The finite element method just presented is nonlinear since $u_{h}^{+}$ is a nonlinear transformation of $u_{h}$. So, it requires an appropriate linearization. In addition, as mentioned earlier, the presence of the term $h_{F}^{-\beta}$ in the penalty can affect the condition number of the linear subproblems tremendously. To address these two issues, we now present a fixed-point iterative algorithm that enables us to both prove the existence of solutions and circumvent the potential ill-conditioning of the scheme. For this, we define the mapping
(11a) $\widetilde{T}\colon V^{0}_{h}\ni w^{0}_{h}\mapsto w^{1}_{h}\mapsto\tilde{w}^{0}_{h}\in V^{0}_{h}\,$,
@@
Step 1: Compute $w^{1}_{h}\in V_{h}^{1}$ by solving the nonlinear scheme: (11b) $a_{h}([P^{w^{0}_{h}}_{h}w_{h}]^{1},v^{1}_{h})+s_{h}([Q^{w^{0}_{h}}_{h}w_{h}]^{1},v^{1}_{h})=(f,v^{1}_{h})_{\mathcal{T}_{h}}-a_{h}(w^{0}_{h},v^{1}_{h})\qquad\forall v^{1}_{h}\in V^{1}_{h}.$ Step 2: Compute $\tilde{w}^{0}_{h}\in V^{0}_{h}$ as solution of (11c) $a_{h}(\tilde{w}^{0}_{h},v^{0}_{h})=(f,v^{0}_{h})_{\mathcal{T}_{h}}-a_{h}(w^{1+}_{h},v^{0}_{h})-\underbrace{s_{h}(w^{1-}_{h},v^{0}_{h})}_{=0}\qquad\forall v^{0}_{h}\in V^{0}_{h}.$
@@
    a function $\tilde{w}_{h}=\tilde{w}_{h}^{1}+\tilde{w}_{h}^{0}$ comprising the fixed-point $\tilde{w}^{0}_{h}$ of $\widetilde{T}$ and the solution of (11b) with $w^{0}_{h}$ replaced by $\tilde{w}^{0}_{h}$ solves (9), which we can easily see by adding (11b) and (11c). So, in the next section we will show that $\widetilde{T}$ has a fixed point;
@@
Remark 3.2 (Bound preservation and mass conservation).
Clearly, $u^{+}_{h}$ takes only values in $\mathcal{G}$ achieving the target described in item 1. Moreover, the stabilization $s_{h}(u_{h}^{-},v_{h})$ vanishes if the $v_{h}\in V_{h}$ is chosen as the characteristic function of one mesh element, i.e., from the space $V^{0}_{h}$. So, the method (9) reduces to the terms only included in (5) for such a test function, and thus the solution $u^{+}_{h}$ satisfies the local mass conservation criterion of item 2 (see also item 3 in Remark 2.1) with the same numerical inter-element flux as the EG scheme defined in (5).
@@ 4 Existence of a discrete solution
This section establishes the existence and stability of a solution to (9), which is made explicit in the following Theorem:
Theorem 4.1.
@@
(12) $\|[\![u^{0}_{h}]\!]\|^{2}_{\mathcal{F}_{h}}\leq\,C\,\frac{h^{2\beta-4}}{\gamma^{2}}\left[C_{P}^{2}\mu^{-1}\|f\|^{2}_{0,\Omega}+(C_{P}^{2}\mu+\epsilon)\max(|a|,|b|)^{2}|\Omega|\right]\,$,
@@
To show this result, we will exploit the fixed-point iteration defined in (11a). Thus, we need to show that the subproblems (11b) and (11c) are well-posed, i.e., that Step 1 and Step 2 in the construction of $\tilde{T}$ are well-defined and continuous operations. We do this in Lemmas 4.12 and 4.14, respectively. Afterward, we establish the fact that $\tilde{T}$ admits at least one fixed-point by applying Brouwer’s fixed point theorem in an adequately constructed ball, cf. Section 4.3. To this end, we use that $\tilde{T}$ is continuous (which follows from Lemmas 4.12 and 4.14), and and show that it maps the adequately constructed ball to itself. We start this endeavor by providing auxiliary results.
@@ 4.1 Preliminaries and auxiliary results
We exploit our fixed-point iteration (11a) to prove this result. Thus, we need to prove that (11b) and (11c) induce well-posed problems, which we perform in Lemmas 4.10 and 4.14, respectively. To this end, we need several auxiliary results.
Lemma 4.2 (Broken Poincaré inequality, [Brenner03]).
@@
Lemma 4.3.
@@
$\left(\mu\|v_{h}^{1}\|_{\mathcal{T}_{h}}^{2}+\epsilon\|\nabla v_{h}^{1}\|_{\mathcal{T}_{h}}^{2}\right)^{1/2}=\|v^{1}_{h}\|_{a_{h}}\leq\frac{C_{\textup{equiv}}}{\sqrt{\alpha}}\|v^{1}_{h}\|_{s_{h}}\qquad\text{ for all }v^{1}_{h}\in V^{1}_{h}.$
Proof 4.4.
The proof of a very similar result can be seen in [BarrenecheaGPV23, Eq. (3.8)], but we summarize it here for completeness. First, using the following inverse inequality (see [EG21-I, Lem. 12.1]): There is a constant $\tilde{c}>0$, independent of $h$ such that
(13) $\tilde{c}h_{T}^{2}|v^{1}_{h}|^{2}_{1,T}\leq\|v^{1}_{h}\|^{2}_{0,T},$
for all $v^{1}_{h}\in V^{1}_{h}$. In addition, we recall the following inequality, whose proof can be found, e.g., in [EG21-I, Prop. 12.5]:
(14) $\|v^{1}_{h}\|^{2}_{0,T}\leq\,C\sum_{\boldsymbol{x}_{i}\in T}h_{i}^{d}\,[v^{1}_{h}(\boldsymbol{x}_{i})]^{2}\,$,
for all $v^{1}_{h}\in V^{1}_{h}$. Then, using the inverse inequality (13) followed by (14) we get to
$\mu\|v_{h}^{1}\|_{\mathcal{T}_{h}}^{2}+\epsilon\|\nabla v_{h}^{1}\|_{\mathcal{T}_{h}}^{2}$ $\leq C\,\sum_{T\in\mathcal{T}_{h}}(\mu+\epsilon h_{T}^{-2})\|v_{h}^{1}\|_{0,T}^{2}$
$\leq C\,\sum_{T\in\mathcal{T}_{h}}(\mu+\epsilon h_{T}^{-2})\sum_{\boldsymbol{x}_{i}\in T}h_{i}^{d}[v^{1}_{h}(\boldsymbol{x}_{i})]^{2}$
$\leq\frac{C}{\alpha}\,s(v_{h}^{1},v_{h}^{1})\,$,
where in the last step we used the mesh regularity. This finishes the proof.
Lemma 4.5.
Given $v^{0}_{h}\in V^{0}_{h}$, there is a constant $C$ (independent of $h$ and $v^{0}_{h}$) such that
$\|v^{1+}_{h}\|^{2}_{0,\Omega}\leq C\left[\max(|a|,|b|)^{2}|\Omega|+\|v^{0}_{h}\|^{2}_{0,\Omega}\right],\quad\text{ for all }v^{1}_{h}\in V^{1}_{h}.$
Proof 4.6.
Let $v_{h}^{1}\in V_{h}^{1}$. The definition of $v^{1+}_{h}(\boldsymbol{x}_{i})$ implies that $|v^{1+}_{h}(\boldsymbol{x}_{i})|\leq\max(|a|,|b|)+\|v^{0}_{h}\|_{\infty,\omega_{i}}$. Using this fact, the inverse inequality (13), the scaling (14), and the mesh regularity we get to
@@
which finishes the proof.
In addition, we also recall the following trace-inverse inequality (see [EG21-I, Lem. 12.10]):
Lemma 4.7.
There exists $C>0$, depending only on the mesh regularity such that for any $T\in{\mathcal{T}_{h}}$ and a facet $F$ of $T$, and any $v_{h}\in V_{h}$ the following holds
(15) $\|v_{h}\|_{0,F}\leq C\,h_{T}^{-1/2}\|v_{h}\|_{0,T}\,.$
The next two results are instrumental in the proof that the nonlinear problem from Step 1 (cf. (11b)) is well-posed. The first result is an generalization of [BarrenecheaGPV23, Lem. 3.1] to the present (more involved) case, so we present a summarized proof.
Lemma 4.8.
@@
Proof 4.9.
Proving the inequality works by going through the below cases. While each of the cases is easy to verify (see also [BarrenecheaGPV23, Lem. 3.1] to this end), it is essential to observe that they cover all possible combinations that $r_{h}^{1}$ and $v^{1}_{h}$ can create:
  • $r_{1},v_{1}\geq b-\overline{w}_{hi}$ or $r_{1},v_{1}\leq a-\underline{w}_{hi}$: In this case $P^{w^{0}_{h}}_{h}r^{1}_{h}]^{1}=[P^{w^{0}_{h}}_{h}v^{1}_{h}]^{1}$.
  • $a-\underline{w}_{hi}\leq r_{1},v_{1}\leq b-\overline{w}_{hi}$: In this case $Q^{w^{0}_{h}}_{h}r^{1}_{h}]^{1}=[Q^{w^{0}_{h}}_{h}v^{1}_{h}]^{1}$.
  • $v_{1}\leq a-\underline{w}_{hi}\leq r_{1}$: $P^{w^{0}_{h}}_{h}v^{1}_{h}]^{1}=a-\underline{w}_{hi}$, $P^{w^{0}_{h}}_{h}r^{1}_{h}]^{1}\geq a-\underline{w}_{hi}$, $Q^{w^{0}_{h}}_{h}v^{1}_{h}]^{1}\leq 0$, $Q^{w^{0}_{h}}_{h}r^{1}_{h}]^{1}\geq 0$.
  • $r_{1}\leq a-\underline{w}_{hi}\leq v_{1}$: $P^{w^{0}_{h}}_{h}r^{1}_{h}]^{1}=a-\underline{w}_{hi}$, $P^{w^{0}_{h}}_{h}v^{1}_{h}]^{1}\geq a-\underline{w}_{hi}$, $Q^{w^{0}_{h}}_{h}r^{1}_{h}]^{1}\leq 0$, $Q^{w^{0}_{h}}_{h}v^{1}_{h}]^{1}\geq 0$.
  • $v_{1}\leq b-\overline{w}_{hi}\leq r_{1}$: $P^{w^{0}_{h}}_{h}v^{1}_{h}]^{1}=b-\overline{w}_{hi}$, $P^{w^{0}_{h}}_{h}r^{1}_{h}]^{1}\leq b-\overline{w}_{hi}$, $Q^{w^{0}_{h}}_{h}v^{1}_{h}]^{1}\geq 0$, $Q^{w^{0}_{h}}_{h}r^{1}_{h}]^{1}\leq 0$.
  • $r_{1}\leq b-\overline{w}_{hi}\leq v_{1}$: $P^{w^{0}_{h}}_{h}r^{1}_{h}]^{1}\leq b-\overline{w}_{hi}$, $P^{w^{0}_{h}}_{h}v^{1}_{h}]^{1}=a-\underline{w}_{hi}$, $Q^{w^{0}_{h}}_{h}r^{1}_{h}]^{1}\leq 0$, $Q^{w^{0}_{h}}_{h}v^{1}_{h}]^{1}\geq 0$.
Lemma 4.10.
@@
[…]
Proof 4.11.
The proof follows exactly the same lines as that of [BarrenecheaGPV23, Thm. 3.2] with [BarrenecheaGPV23, Lem. 3.1: (25)] replaced by Lemma 4.8.
@@ 4.2 Well-posedness of the iteration defined in (11a)
Lemma 4.12 (Well-posedness of (11b)).
Under the assumptions of Lemma 4.10, the equation (11b) uniquely defines $w^{1}_{h}\in V^{1}_{h}$ for a given $w^{0}_{h}$. Moreover, if $\hat{w}^{0}_{h},\bar{w}^{0}_{h}\in V^{0}_{h}$, and $\hat{w}^{1}_{h},\bar{w}^{1}_{h}\in V^{1}_{h}$ are their images under $\tilde{T}_{1}$, then the following Lipschitz continuity holds
@@
Proof 4.13.
The existence and uniqueness of the solution of (11b) are a direct consequence of the strong monotonicity and continuity of $\tilde{T}_{1}$ (see, e.g., [RenardyR04, Thm. 10.49]). To prove the Lipschitz continuity, let $\hat{w}_{h}^{0},\bar{w}_{h}^{0}\in V_{h}^{0}$ and $\hat{w}_{h}^{1},\bar{w}_{h}^{1}\in V_{h}^{1}$ be their images under $\tilde{T}_{1}$. Then, using (17) and the problem (11b) we get to
@@
and the proof is finished using the continuity of $a_{h}(\cdot,\cdot)$.
@@
Lemma 4.14 (Well-posedness of (11c)).
@@
(18) $a_{h}(w^{0}_{h},v^{0}_{h})=\mu(w^{0}_{h},v^{0}_{h})_{\mathcal{T}_{h}}+\gamma\left\langle\frac{\epsilon+\mu h_{F}^{2}}{h_{F}^{\beta}}[\![w^{0}_{h}]\!],[\![v^{0}_{h}]\!]\right\rangle_{\mathcal{F}_{h}}\qquad\text{ for all }w^{0}_{h},v^{0}_{h}\in V^{0}_{h}.$
@@
for every $v_{h}^{0}\in V_{h}^{0}$. Taking $v_{h}^{0}=\hat{w}^{0}_{h}-\bar{w}^{0}_{h}$ and using Cauchy-Schwarz’s inequality on the right-hand side above leads to the result.
4.3 Proof of Theorem 4.1
@@
Proof 4.17.
Let $v^{0}_{h}=\tilde{w}^{0}_{h}$ in (11c) and $\tilde{\gamma}=\tfrac{\gamma}{h^{\beta-1}}$. Then, using (18), the ellipticity of $a_{h}(\cdot,\cdot)$ in $V_{h}^{0}$, (11c), Cauchy-Schwarz’s inequality, the local trace-inverse inequality (15), the broken Poincaré inequality from Lemma 4.2, and a global inverse inequality, we get to
$\|\sqrt{\mu}\tilde{w}^{0}_{h}\|^{2}_{0,\Omega}+\tilde{\gamma}\|[\![\tilde{w}^{0}_{h}]\!]\|^{2}_{\mathcal{F}_{h}}=a_{h}(\tilde{w}^{0}_{h},\tilde{w}^{0}_{h})$
@@
The proof then follows from rearranging terms.
Lemma 4.18.
Let us assume that $\beta\geq 2$. Then, if $\gamma>0$ is sufficiently large, $\tilde{T}\colon B_{1}(0)\to B_{1}(0)$, where $B_{1}(0)\subset V^{0}_{h}$ is the closed unit ball in $(V_{h}^{0},\|[\![\cdot]\!]\|_{\mathcal{F}_{h}})$.
Proof 4.19.
Using Lemma 4.5, Lemma 4.16, and the broken Poincaré inequality from Lemma 4.2 we get to
@@
(19) $\leq\,C\,\frac{h^{2\beta-4}}{\gamma^{2}}\left[C_{P}^{2}\mu^{-1}\|f\|^{2}_{0,\Omega}+(C_{P}^{2}\mu+\epsilon)\max(|a|,|b|)^{2}|\Omega|+C_{P}^{2}\frac{C_{p}^{2}\mu+\epsilon}{\epsilon+\mu h^{2}}\|[\![w^{0}_{h}]\!]\|^{2}_{{\mathcal{F}_{h}}}\right]\,.$
The result is then proved by noticing that $\gamma$ can always be chosen large enough so that the last term in the right-hand side above can be hidden in the left-hand side, and the resulting right-hand side is smaller than $1$.
The proof of Theorem  4.1 appears as a corollary of the above results.
Proof 4.20 (Proof of Theorem 4.1).
Since the EG finite element space is finite-dimensional, the operator $\tilde{T}$ is continuous and maps the closed unit ball (a convex compact set) onto itself. Thus, Brouwer’s fixed-point Theorem (see, e.g. [RenardyR04, Thm. 10.41]) states that $\tilde{T}$ has at least one fixed point in $B_{1}(0)$, which is the desired result.